# Microbiomes Through the Looking Glass

### v1

@@ 2.3 Species interactions patterns characterize the state of microbiomes
We thus collect all the parameters estimated from the data in a vector $\boldsymbol{\pi}=(h,q_{d},q_{0},K)$. As we will better detail in the Methods, we develop a moment matching inference algorithm to infer the model parameters $\boldsymbol{\theta}$. We therefore generated order parameters from our model that are as close as possible to the data, as depicted in Figure 1. The idea of the method is to introduce a cost function $\mathcal{C}(\boldsymbol{\theta}|\boldsymbol{\pi})$, representing a total relative error for some self-consistent equations. If the parameters $\boldsymbol{\theta}$ are such that the right part of the self-consistent equation equals the left part, the problem can be considered solved. Since the landscape of this cost function presents several minima, we performed multiple optimization procedures to collect an ensemble of possible solutions, from which we retained the top 30.
@@
First, we find that different solutions $\boldsymbol{\theta}^{*}$ of the optimization problem provides ecological insights into the underlying microbiome populations. As Figure 3A shows, the inferred demographic noise strength ($T$) and the heterogeneity of the interactions ($\sigma$) for healthy and diseased microbiomes are clustered in the $T-\sigma$ plane. In particular, the SAD for the healthy cohort is robust among the different solutions of the inference procedure, as shown by the superposition of the different curves in the inset of Figure 3A. On the other hand, SADs inferred from unhealthy patients have high sensitivity to different solutions. In particular, some of them display a mode for high-abundance species (light red lines in Figure 3A), a signature of dominant strain in the gut. Consistently, the distribution of the interactions $P(\alpha_{i,j})$ generated through the inferred parameters $\mu$ and $\sigma$ is different between healthy and diseased cohorts, giving a distinct pattern of interactions (see Figure 3B), a result that is compatible with that found by Bashan and collaborators Bashan et al. (2016). In particular, we find that dysbiosis reduces the heterogeneity of interaction strengths, a result also observed when taking correlations as a proxy for interactions Seppi et al. (2023).
Figure 3: Panel a: inferred $T$ (demographic noise strength) and $\sigma$ (interactions heterogeneity) for healthy (blue) and diseased (red) microbiomes are clustered. Darker dots correspond to better solutions (i.e., solutions with a lower value of the cost function $\mathcal{C}$), while the two points with hexagonal markers correspond to the best two (healthy and diseased, respectively) solutions. In the first panel inset, we also show (in log-log scale) the SADs corresponding to each solution. To have a more concise representation, we present each SAD fixing the disorder to its average $\overline{\zeta}=K-\mu h$. Panel b: the probability density function of the inferred interactions $\alpha_{i,j}$ for healthy (blue) and diseased (red) microbiomes. Dysbiosis reduces the heterogeneity of the interaction strengths.
@@
We then assess how close the inferred $\sigma$ and $\beta=1/T$ (the inverse temperature in a statistical physics approach) are to the critical replica symmetry-breaking (RSB) line of the dgLV ($\mathcal{R}=0$), evaluated being all the other parameters constant (see Methods). We find again that the replicon values $\mathcal{R}$ corresponding to each solution of our optimization protocol are significantly different for the two investigated microbiome phenotypes (see Figure 4A). In particular, diseased microbiomes are closer to marginal stability within the replica-symmetric ansatz Altieri et al. (2021); Mézard et al. (1987); de Almeida and Thouless (1978). Furthermore, by investigating the shape of the SAD given by Eq. (2), we can estimate the ratio between niche (represented by species interaction) and neutral (represented by birth/death and immigration) ecological forces, which can be captured by the quantity $\psi$ Wu et al. (2021). It detects the emergence of internal modes in the species abundance distribution as a hallmark of niche processes (see SI, Section E).
@@
Inspired by field-theory arguments (see Methods section Free-energy landscape exploration: replica formulation), we can call mass of the theory the $m$ parameter as defined above in Eq. (2). When $m\rightarrow 0$ the theoretical order parameters diverge and the system enters a pathological regime of unbounded growth. As such, the mass term, can be considered as a complementary stability measure, capable of capturing the transition to the unbounded growth regime.
@@
In the model, two kinds of effects compete to shape the community structure. On the one hand, we have niche effects, encoded in disordered interactions and thus tracked by the parameters $\mu$, $\sigma$, and $K$. Their overall effect is selective and tends to concentrate the SAD around the typical abundance value. On the other hand, we have neutral effects encoded into the stochastic dynamics and immigration, governing the low abundance regime of the SAD. When the demographic noise amplitude is stronger than immigration ($\nu<1$, as in our case) the SAD exhibits a low-abundance integrable divergence. In the opposite scenario, $\nu>1$, there is no divergence and the SAD is modal. Since interactions are random, the probability of observing an internal mode can be estimated as the fraction of SADs realization having non-trivial solutions to the stationary point equation. Such quantity, dubbed as the niche-neutral ratio, can be analytically evaluated and reads:
@@
where $\zeta^{*}=\sqrt{\frac{4(1-\nu)m}{\beta}}$ and $\overline{\zeta}=K-\mu h$. When $\psi\approx 1$, niche and neutral forces give comparable contributions to the dynamics, as both low abundance divergence and a finite abundance mode coexist in the SAD. Finally, if the typical abundance diverges, we enter into the unbounded growth phase, which means that the mass $m$ and the niche-neutral ratio $\psi$ are not independent, as suggested by the analytical expression on $\psi$. For an exhaustive derivation of this result, see SI (Section E). With the model parameters obtained, we were able to evaluate $m$ and $\psi$ for healthy and diseased microbiomes. Also, in this case, healthy and diseased microbiomes are clearly clustered, as shown in Figure (4). Unhealthy microbiomes turn out to be closer to the unbounded growth phase, and the nice-neutral ratio is larger by five orders of magnitude compared to the healthy case $\psi_{U}\approx 10^{5}\psi_{H}$. This leads us to argue that selective pressure is way larger in diseased states, while in the healthy one, birth and death effects are the key drivers of the dynamics. These results are also confirmed by the SADs shape in the inset of Figure (3, panel a).
@@
Figure 4: The replicon eigenvalue corresponding to each solution of our optimization procedure (shaded dots). The solid hexagon represents the replicon corresponding to the best solutions that minimize the error in predicting the order parameters of the theory (minimum $\mathcal{C}$). The two investigated microbiome phenotypes (healthy in blue, diseased in red) are significantly different. In particular, diseased microbiomes are closer to the marginal stability of replica-symmetric ansatz (grey horizontal line). Panel b: Solutions of the moment-matching objective function are shown as a function of $\psi$ and $m$, which in turn depend on the SAD parameters (see main text). Healthy (blue) and diseased (red) microbiomes appear to be clustered. Therefore, distinct ecological organization scenarios (strong neutrality/emergent neutrality) emerge. Darker dots correspond to solutions with lower values of the cost function, while hexagonal markers correspond to the two best solutions.
Discussion and Conclusions
@@
In our exploration of the gut microbiome through the lens of disordered systems and random matrix theory, we have embarked on an ambitious journey to connect intricate theoretical concepts to practical analyses of environmental microbiome data. In particular, we have characterized healthy and unhealthy gut microbiomes using the disordered generalized Lotka-Volterra model (dgLV) for population dynamics. This attempt to apply advanced theoretical methodologies from disordered systems to real-world biological data represents a significant stride in understanding the complex interplay within microbial communities.
One delicate aspect of our study lies in addressing the inherent limitations and the potential scope for improvement in future research. A notable challenge is the discrepancy between the empirical Species Abundance Distribution (SAD), which is observed in the data, and the theoretical SAD predicted by the quenched dgLV model Suweis et al. (2023). While empirical data showcase a diverse range of species abundance, following a power-law distribution, the model predictions tend to exhibit exponential decay in SAD tails. This mismatch underscores the need for further refinement of the model to accurately capture the nuanced patterns observed in real-world data. For example, recently it has been proposed that introducing an annealed disorder can generate SADs that look closer to the empirical ones Suweis et al. (2023). Another possibility is to set the dgLV model parameters in such a way as to reproduce the multi-attractor phase. In fact, in this region, the SADs display a more heterogeneous shape Arnoulx de Pirey and Bunin (2024). We already intend to explore this follow-up direction by combining one-step replica-symmetry-breaking (1RSB) computations in the multiple-attractor phase with Dynamical Mean-Field Theory analysis for the asymmetric interaction case.
@@
Another related limitation is the challenge of generating species abundance samples from the dgLV model that mirrors the statistical properties of the observed empirical data. In our current framework, each microbiome sample could be extracted from $p(N|\zeta)$, where $\zeta$ is a realization of the disorder. However, $p(N|\zeta)$ near $N\rightarrow 0$ presents a power-law exponent $\nu-1$, with $\nu_{H,D}\approx 10^{-3}$. This generates numerical instabilities and dominates the sampling process, posing difficulties in generating representative synthetic samples. Moreover, while microbiome data are inherently compositional Pasqualini et al. (2023), the dgLV model species populations $N_{i}$ are positive real numbers. However, as already noted, it can be shown that introducing a normalization $\tilde{N}_{i}\rightarrow N_{i}/\sum_{j}N_{j}$ in such equations does not change the structure of the proposed solutions Altieri et al. (2021) and therefore should not affect the conclusions of our work. The study also sheds light on the role of demographic noise within the context of the dgLV model. In fact, in the limit $T\rightarrow 0$, the SAD transitions to a truncated Gaussian, with a prominent peak at zero reflecting the fraction of extinct species. In contrast to scenarios at zero noise, where species extinctions are observed, the inclusion of demographic noise in the dgLV model suggests a scenario where no species goes extinct, supporting the everything is everywhere hypothesis Grilli (2020); Pigani et al. (2024). In other words, within this framework, zeros in the data are due to sampling effects and not to local species extinctions Grilli (2020); Pasqualini et al. (2023).
@@
In conclusion, our work opens up new avenues for future research, particularly in refining the theoretical models to better align with empirical observations and in exploring the nuances of species abundance distributions within the microbiome context. Moreover, the integration of other forms of environmental variability and species-specific traits could provide a more holistic view of ecosystem dynamics as also proposed by the One Health-One Microbiome framework Tomasulo et al. (2024). The insights gained from this study not only enhance our understanding of the complex dynamics in microbial ecosystems but also pave the way for more sophisticated analytical tools in ecological modelling, contributing significantly to our ability to interpret and manage these vital biological systems.
Aknowledgements

### v2

@@ 2.3 Species interactions patterns characterize the state of microbiomes
We thus collect all the parameters estimated from the data in a vector $\boldsymbol{\pi}=(h,q_{d},q_{0},K)$. As we will better detail in the Methods, we develop a moment matching inference algorithm to infer the model parameters $\boldsymbol{\theta}$, as depicted in Figure 1. The idea of the method is to introduce a cost function $\mathcal{C}(\boldsymbol{\theta}|\boldsymbol{\pi})$, representing a total relative error for some self-consistent equations. If the parameters $\boldsymbol{\theta}$ are such that the right part of the self-consistent equation equals the left part, the problem is considered solved. Because the landscape associated with this cost function presents several minima, we perform multiple optimization procedures to collect an ensemble of possible solutions, from which we retain the top 30. First, we find that different solutions $\boldsymbol{\theta}^{*}$ of the optimization problem provide ecological insights into the underlying microbiome populations.
As originally predicted in Altieri et al. (2021), among all the parameters that define Eq. (1), the only ones relevant for reproducing the theoretical phase diagram are the amplitude of demographic noise and the heterogeneity of interactions. The mean interaction strength, provided it is sufficiently positive, does not play a significant role. This prediction is fully confirmed by the inference procedure applied to the two microbiome datasets, allowing us to identify a universal signature that distinguishes healthy from unhealthy states. Figure 3a shows, indeed, that inferred noise ($T$) and interaction heterogeneity strength ($\sigma$) for healthy and diseased microbiomes are clustered in the two-dimensional plane.
@@
In particular, the SAD for the healthy cohort is robust among the different solutions of the inference procedure, as depicted by the superposition of the different curves in the inset of Figure 3A. On the other hand, SADs inferred from unhealthy patients have high sensitivity to different solutions. In particular, some of them display a mode for high-abundance species (light red lines in Figure 3a), a signature of dominant strain in the gut. Consistently, the distribution of the interactions $P(\alpha_{i,j})$ generated through the inferred parameters $\mu$ and $\sigma$ is different between healthy and diseased cohorts, giving a distinct pattern of interactions (see Figure 3b), a result that is compatible with that found by Bashan and collaborators Bashan et al. (2016). Remarkably, we find that dysbiosis reduces the heterogeneity of interaction strengths, a result also observed when taking correlations as a proxy for interactions Seppi et al. (2023).
@@
We then assess how close the inferred $\sigma$ and $\beta=1/T$ (a.k.a. inverse temperature in a statistical physics approach) are to the critical replica symmetry-breaking (RSB) line of the dgLV ($\mathcal{R}=0$), evaluated by keeping all the other parameters constant (see Methods). We find again that the replicon values $\mathcal{R}$ corresponding to each solution of our optimization protocol are significantly different for the two investigated microbiome phenotypes (see Figure 4a). In particular, diseased microbiomes are closer to marginal stability within the replica-symmetric ansatz Altieri et al. (2021); Mézard et al. (1987); de Almeida and Thouless (1978). Furthermore, by investigating the shape of the SAD given by Eq. (2), we can estimate the ratio between niche (represented by species interaction) and neutral (represented by birth/death and immigration) ecological forces, which can be captured by the quantity $\psi$ Wu et al. (2021). It detects the emergence of peaks in the species abundance distribution as a hallmark of niche processes (see Appendix 2).
@@
Inspired by field-theory arguments (see Methods and Appendix, Section S2), we can call mass of the theory the $m$ parameter, as defined above in Eq. (2). In classical and quantum field theory, the particle-particle interaction embedded in the quadratic term is typically referred to as a mass source. In our context, $m=1-\beta\sigma^{2}(q_{d}-q_{0})$ captures quadratic fluctuations of species abundances, as also appearing in the expression of the leading eigenvalue of the stability matrix. When $m\rightarrow 0$, the analytical order parameters diverge and the system enters the unphysical regime of unbounded growth. As such, the mass term can be considered a complementary stability measure, capable of capturing the transition to the unbounded growth regime.
Figure 3: Panel a: inferred $T$ (demographic noise strength) and $\sigma$ (interactions heterogeneity) for healthy (blue) and diseased (red) microbiomes are clustered. Darker dots correspond to better solutions (i.e., solutions with a lower value of the cost function $\mathcal{C}$), while the two points with hexagonal markers correspond to the best two (healthy and diseased, respectively) solutions. In the first panel inset, we also show (in log-log scale) the SADs corresponding to each solution. To have a more concise representation, we present each SAD fixing the disorder to its average $\overline{\zeta}=K-\mu h$. Panel b: the probability density function of the inferred interactions $\alpha_{i,j}$ for healthy (blue) and diseased (red) microbiomes. Dysbiosis reduces the heterogeneity of the interaction strengths. The quantities reported in the legend are the average and standard deviation of $\alpha_{i,j}$. They have are calculated as $\mu^{\alpha}_{X}=\mu_{X}/S_{X}$ and $\sigma^{\alpha}_{X}=\sigma_{X}/\sqrt{S_{X}}$, where $S_{X}$ is the species pool size, estimates as the set of all the observed species in a dataset, $X$. $X$ can be healthy ($H$) or diseased ($D$).
@@
In the model, two kinds of effects compete to shape the community structure. On the one hand, we have niche effects, encoded in disordered interactions and thus tracked by the parameters $\mu$, $\sigma$, and $K$. Their overall effect is selective and tends to concentrate the SAD around the typical abundance value. On the other hand, we have neutral effects encoded in the stochastic dynamics and immigration, governing the low abundance regime of the SAD. When the demographic noise amplitude is stronger than immigration ($\nu<1$, as in our case), the SAD exhibits a low-abundance integrable divergence. In the opposite scenario, for $\nu>1$, there is no divergence and the SAD is modal. Since interactions are random, the probability of observing an internal mode can be estimated as the fraction of SADs realizations having non-trivial solutions to the stationary point equation. Such a quantity, dubbed as the niche-neutral ratio, can be analytically evaluated:
@@
where $\zeta^{*}=\sqrt{\frac{4(1-\nu)m}{\beta}}$ and $\overline{\zeta}=K-\mu h$. When $\psi\approx 1$, niche and neutral forces give comparable contributions to the dynamics, as both low abundance divergence and a finite abundance mode coexist in the SAD. Finally, if the typical abundance diverges, we enter the unbounded growth phase, which means that the mass $m$ and the niche-neutral ratio $\psi$ are not independent, as suggested by the analytical expression for $\psi$. For an exhaustive derivation of this result, see Appendix 2. With the obtained model parameters, we are able to evaluate $m$ and $\psi$ for healthy and diseased microbiomes. Also in this case healthy and diseased microbiomes are visibly clustered, as shown in Figure 4. Unhealthy microbiomes turn out to be closer to the unbounded growth phase, and the niche-neutral ratio is larger by five orders of magnitude than the healthy case $\psi_{D}\approx 10^{5}\psi_{H}$. This leads us to argue that selective pressure is way larger in diseased states, while in the healthy one, birth and death effects are the key drivers of the dynamics. These results are also confirmed by the SAD shapes in the inset of Figure 3 (panel a).
@@
Figure 4: Panel a: The replicon eigenvalue corresponding to each solution of our optimization procedure (shaded dots). The solid hexagon represents the replicon corresponding to the best solutions that minimize the error in predicting the order parameters of the theory (minimum $\mathcal{C}$). The two investigated microbiome phenotypes (healthy in blue, diseased in red) are significantly different. In particular, diseased microbiomes are closer to the marginal stability of replica-symmetric ansatz (grey horizontal line). Panel b: Solutions of the moment-matching objective function are shown as a function of $\psi$ and $m$, which in turn depend on the SAD parameters (see main text). Healthy (blue) and diseased (red) microbiomes appear to be clustered. Therefore, distinct ecological organization scenarios (strong neutrality/emergent neutrality) emerge. Darker dots correspond to solutions with lower values of the cost function, while hexagonal markers correspond to the two best solutions.
In summary, in the Result section we demonstrate that: (i) the inference pipeline robustly recovers demographic noise and interaction heterogeneity by calculating $h$, $q_{0}$ and $q_{d}$ from the data; (ii) These parameters cluster by health status and (iii) diseased microbiomes lie closer to the replica-symmetry-breaking threshold, indicating reduced ecological resilience.
3 Discussion and Conclusions
@@
In our exploration of the gut microbiome through the lens of disordered systems and random matrix theory, we have proposed a connection between the theoretical framework of disordered systems and practical analyses of environmental microbiome data. In particular, we have characterized healthy and unhealthy gut microbiomes using the disordered generalized Lotka-Volterra model (dgLV) for population dynamics. We now interpret our theoretical findings in a biological context, contrast them with previous work, and outline limitations and future directions.
The first major result of our work suggests a different role for the various ecological forces shaping the human gut microbial community. In this sense, the niche-neutral ratio $\psi$, highlights the different roles of interactions in healthy and diseased microbiomes. In the healthy case, neutral forces, such as random birth and death of individuals, characterize the dynamics, making configurations corresponding to this state alike. On the contrary, in diseased microbiomes, disordered, sample-specific interactions are the dominant ecological force, making individual realizations differ significantly from one another. An ecological interpretation of our findings suggests that healthy microbiomes are governed primarily by demographic stochasticity, reflecting a quasi-neutral regime characterized by similar community structures across individuals. Conversely, microbiomes from diseased patients exhibit significantly greater variability, suggesting that deterministic ecological factors – such as weakened species interactions – override neutrality, leading to structural instability and distinct microbial compositions. This observation aligns with the "Anna Karenina principle" Pasqualini et al. (2024); Ma (2020) holding for gut microbiomes, which can be phrased as: "All healthy gut microbiomes are alike; each unhealthy gut microbiome is unhealthy in its own way”. Supporting this interpretation, our analysis of the replicon eigenvalue $\mathcal{R}$ shows that the healthy state is associated with pronounced stability to external perturbations. The unhealthy state, instead, being closer to the RSB line, exhibits diminished stability and, consequently, reduced robustness against external perturbations.
@@
Our study also sheds light on the role of demographic noise within the context of the dgLV model. In the limit $T\rightarrow 0$, the SAD transitions to a truncated Gaussian, with a prominent peak at zero reflecting the fraction of extinct species. In contrast to a zero-noise scenario, where species extinctions are observed, the inclusion of demographic noise in the dgLV model suggests a picture where no species goes extinct, supporting the everything is everywhere hypothesis Grilli (2020); Pigani et al. (2024). In other words, within this framework, zeros in the data are due to sampling effects and not due to local species extinctions Grilli (2020); Pasqualini et al. (2024).
A notable limitation of our study lies in the discrepancy between the empirical Species Abundance Distribution (SAD) observed in the data and the theoretical distribution predicted by the quenched dgLV model Suweis et al. (2024). While empirical data showcase a diverse range of species abundance, following a power-law distribution, the model predictions tend to exhibit exponential decay in SAD tails. This mismatch underscores the need for further refinement of the model to accurately capture the nuanced patterns observed in real-world data. For example, it has recently been proposed that introducing an annealed disorder [^3] Unlike the quenched approximation, the annealed version assumes that random couplings are not fixed but rather fluctuate over time, with their covariance governed by independent Ornstein–Uhlenbeck processes. can generate SADs that more closely resemble the empirical ones Suweis et al. (2024). Another possibility is to set the dgLV model parameters in such a way as to reproduce the multi-attractor phase. In fact, in this region, the SADs display a more heterogeneous shape Arnoulx de Pirey and Bunin (2024). We plan to explore this follow-up direction by combining one-step replica-symmetry-breaking (1RSB) computations in the multiple-attractor phase with Dynamical Mean-Field Theory analysis for the asymmetric interaction case. This latter approach is particularly well suited for studying inherently non-equilibrium dynamics and for extending the framework to systems subject to environmental fluctuations in addition to demographic noise.
Another related limitation is the challenge of generating species abundance samples from the dgLV model that mirrors the statistical properties of the observed empirical data. In our current framework, each microbiome sample could be extracted from $p(N|\zeta)$, where $\zeta$ is a realization of the disorder. However, $p(N|\zeta)$ near $N\rightarrow 0$ presents a power-law exponent $\nu-1$, with $\nu_{H,D}\approx 10^{-3}$. This results in numerical instabilities and dominates the sampling process, posing difficulties in generating representative synthetic samples. Moreover, while microbiome data are inherently compositional Pasqualini et al. (2024), the dgLV model species populations $N_{i}$ are positive real numbers. However, as already noted, it can be shown that introducing a normalization $\tilde{N}_{i}\rightarrow N_{i}/\sum_{j}N_{j}$ in such equations does not change the structure of the proposed solutions Altieri et al. (2021) and therefore should not affect the conclusions of our work.
@@
In conclusion, our work proposes a bridge between theory and data, particularly in refining the theoretical models to better align with empirical observations and in exploring the nuances of species abundance distributions within the microbiome context. Moreover, the integration of other forms of environmental variability and species-specific traits could provide a more holistic view of ecological dynamics, as also proposed by the One Health-One Microbiome framework Tomasulo et al. (2024).
Overall, our study builds a quantitative link between metagenomic data and the disordered gLV framework, revealing how dysbiosis alters gut species interaction networks. By doing so, it lays the groundwork for more advanced, mechanistically informed models to better interpret and ultimately manage complex microbial ecosystems.
4 Acknowledgments